# Concept Guided Co-saliency Objection Detection

### v1

@@ 5 Experiment
GCAGC [48] 0.8029 0.7162 0.0813 0.8281 0.8888 0.8346
GICD [51] 0.8382 0.7484 0.0744 0.8261 0.8774 0.8314
GCoNet [9] 0.8272 0.7472 0.0708 0.8327 0.8808 0.8324
DCFM [42] 0.8486 0.7384 0.0678 0.8446 0.8883 0.8291
CADC [50] 0.8625 0.7723 0.0687 0.8550 0.8992 0.8591
ConceptCoSOD 0.8719 0.7556 0.0570 0.8292 0.8940 0.8385
@@
GCAGC [48] 0.7026 0.6260 0.0929 0.7556 0.8518 0.7915
GICD [51] 0.7361 0.6564 0.0883 0.7461 0.8326 0.7779
GCoNet [9] 0.7322 0.6584 0.0773 0.7569 0.8464 0.7847
DCFM [42] 0.7677 0.6706 0.0712 0.7853 0.8645 0.7949
CADC [50] 0.7457 0.6737 0.0928 0.7644 0.8465 0.8052
ConceptCoSOD 0.8072 0.7003 0.0700 0.7853 0.8681 0.8070
@@
GCAGC [48] 0.3915 0.3766 0.1082 0.5086 0.7544 0.6684
GICD [51] 0.3683 0.4031 0.1455 0.4757 0.6778 0.6311
GCoNet [9] 0.3984 0.4206 0.1147 0.5089 0.7306 0.6517
DCFM [42] 0.4826 0.4697 0.0901 0.5697 0.7646 0.6914
CADC [50] 0.4239 0.4406 0.1339 0.5326 0.7370 0.6720
ConceptCoSOD 0.5644 0.5184 0.0956 0.5956 0.7593 0.7109
@@
$E^{max}_{\xi}$$\uparrow$ 0.7243 0.7496 0.7593 0.7441 0.7112 0.6390
$S_{m}$$\uparrow$ 0.6669 0.6952 0.7109 0.7090 0.6912 0.6522
Table 3: Co-saliency detection performance on corrupted datasets.
Frost Fog
[…]
GCAGC [48] 0.6714 0.6114 0.1121 0.7548 0.8394 0.7769 0.7548 0.6725 0.0944 0.7944 0.8628 0.8089
GICD [51] 0.7315 0.6589 0.1008 0.7581 0.8256 0.7745 0.7736 0.6892 0.0935 0.7799 0.8443 0.7941
GCoNet [9] 0.6823 0.6196 0.1052 0.7380 0.8049 0.7539 0.7399 0.6697 0.0908 0.7729 0.8357 0.7865
DCFM [42] 0.6848 0.6093 0.1037 0.7462 0.8056 0.7482 0.7548 0.6618 0.0872 0.7879 0.8427 0.7826
CADC [50] 0.7429 0.6683 0.0956 0.7798 0.8469 0.8014 0.8029 0.7205 0.0822 0.8173 0.8734 0.8320
ConceptCoSOD 0.8074 0.7034 0.0837 0.7824 0.8524 0.7983 0.8367 0.7179 0.0760 0.7925 0.8676 0.8089
Motion Blur Defocus Blur
[…]
GCAGC [48] 0.6853 0.6113 0.1240 0.7394 0.8191 0.7700 0.7057 0.6269 0.1246 0.7464 0.8197 0.7747
GICD [51] 0.7052 0.6332 0.1088 0.7290 0.8117 0.7567 0.7126 0.6310 0.1059 0.7317 0.8090 0.7571
GCoNet [9] 0.6203 0.5678 0.1105 0.6935 0.7683 0.7247 0.5781 0.5320 0.1180 0.6578 0.7320 0.7038
DCFM [42] 0.6655 0.5948 0.0997 0.7287 0.7969 0.7404 0.6684 0.5958 0.0995 0.7352 0.7968 0.7422
CADC [50] 0.7260 0.6504 0.1026 0.7538 0.8345 0.7898 0.6928 0.6162 0.1054 0.7310 0.8252 0.7736
ConceptCoSOD 0.7895 0.6772 0.0868 0.7542 0.8403 0.7830 0.7464 0.6492 0.0939 0.7233 0.8087 0.7657
Adversarial Attack (Jadena) Gaussian Noise
[…]
GCAGC [48] 0.5980 0.5501 0.1444 0.6900 0.7837 0.7347 0.5320 0.5156 0.2265 0.6332 0.6912 0.6746
GICD [51] 0.6794 0.6162 0.1110 0.7111 0.7905 0.7481 0.6853 0.6288 0.1224 0.7200 0.7995 0.7470
GCoNet [9] 0.5627 0.5257 0.1178 0.6447 0.7218 0.7014 0.6054 0.5571 0.1183 0.6887 0.7663 0.7155
DCFM [42] 0.5995 0.5404 0.1097 0.6678 0.7412 0.7112 0.6312 0.5659 0.1112 0.7044 0.7760 0.7217
CADC [50] 0.7176 0.6417 0.0978 0.7553 0.8334 0.7886 0.7141 0.6457 0.1070 0.7553 0.8295 0.7852
ConceptCoSOD 0.8004 0.6947 0.0921 0.7722 0.8413 0.7907 0.7756 0.6801 0.0897 0.7584 0.8282 0.7858
5.1 Experimental Setup
Datasets. Our experiments on the clean datasets were conducted on Cosal2015 [45], CoSOD3k [8], and CoCA [52]. These three datasets contain 2015, 3316, and 1295 images of 50, 160, and 80 groups respectively. In terms of the corrupted datasets, we applied six types of corruption methods to the Cosal2015 dataset, specifically including frost, fog, motion blur, defocus blur, and Gaussian noise from common corruptions types [15], as well as the adversarial attack Jadena [13]. Specifically, we applied corruption to the first $50\%$ of images in each image group while keeping the remaining $50\%$ of images clean.
Baseline methods. We selected five comparison methods that have demonstrated excellent performance in co-saliency object detection, including GCAGC [48], GICD [51], GCoNet [9], DCFM [42] and CADC [50].
Implementation details. In the group-image concept learning module, the timestep $t$ is set to 1,000. The learning rate for updating $\textbf{c}^{*}$ is set to $5e^{-4}$, and the max train step is set to 2,000. The threshold for segmentation is set to 0.5.
5.2 Comparison on Clean Datasets
We compare our ConceptCoSOD method with baselines on Cosal2015, CoSOD3k, and CoCA datasets. As shown in Table 1, the columns represent the baseline models and our method. The top two values for each metric are highlighted, with red indicating rank-1 and yellow indicating rank-2. For the Cosal2015 dataset, our method achieves five top-2 rankings, matching CADC’s performance. On the CoSOD3k dataset, our method attains rank-1 across all six metrics. For the CoCA dataset, our method secures six top-2 rankings, equivalent to DCFM. Overall, our method demonstrates the best performance across all three datasets.
5.3 Comparison on Corrupted Datasets
The robustness of our co-salient object detection method is essential, as real-world images often include various corruptions. To assess robustness, we applied common corruptions and an adversarial attack, as presented in Table 3. Specifically, we evaluated weather-related corruptions (frost, fog), common distortions (motion blur, defocus blur, Gaussian noise) [15], and an adversarial attack (Jadena) [13]. The top two scores for each metric are highlighted, with red indicating the highest and yellow indicating the second highest. Results show that our method achieves top performance under frost, motion blur, adversarial attacks, Gaussian noise, and defocus blur. For fog corruption, our method performs slightly below CADC but remains highly competitive. Overall, our method demonstrates superior robustness across a variety of corruptions.
5.4 Visualization
As shown in Fig. 7, we demonstrate the visualization result comparison between our method and baselines under corruptions. There are three cases, each containing six types of corruption. In most instances, only our ConceptCoSOD method effectively segments the target object, while other methods largely fail to identify it. This outcome demonstrates the superior robustness of our method against corruption compared to others.
Figure 8: Ablation on concept guidance.
5.5 Ablation Study
Concept guidance. We utilize different concepts (text embeddings) to assess their influence on segmentation performance. As shown in Fig. 8, the input image is of the dog concept. We observe that segmentation using the concept learned from the dog image group achieves better results in segmenting the dog object compared to the concept learned from the mouse image group. Also, if we directly put the text embedding of the word “dog” as guidance, the segmentation result is good, reflecting the effect of text semantics as supplementary information. If we do not put any text as guidance, the result is black, which is totally wrong.
@@ 5.6 Limitation
According to the robustness analysis, our method does not yet outperform other Co-SOD approaches comprehensively. In future work, we aim to enhance the robustness through advanced data augmentation techniques.
@@ 6 Conclusion
This paper presents a novel approach, ConceptCoSOD, which leverages text-based semantic guidance for co-salient object detection (Co-SOD). Unlike traditional Co-SOD methods that rely solely on image-based features, ConceptCoSOD introduces a robust (image-text)-to-image framework, enhancing object segmentation accuracy by incorporating semantic information from text. This approach not only mitigates the challenges posed by inter-object variations and background noise but also showcases the potential of integrating text-driven semantics into Co-SOD tasks, providing a new direction for future research in vision-language models for complex segmentation tasks. In future work, we aim to enhance the robustness of concept learning to improve the method’s practicality.

### v2

@@ 5 Experiment
5.1 Experimental Setup
Datasets. Our experiments on the clean datasets were conducted on Cosal2015 [44], CoSOD3k [8], and CoCA [50]. These three datasets contain 2015, 3316, and 1295 images of 50, 160, and 80 groups respectively. In terms of the corrupted datasets, we applied five types of corruption methods to the Cosal2015 dataset, specifically including frost, motion blur, defocus blur, and Gaussian noise from common corruption types [16], as well as the adversarial attack Jadena [14]. Specifically, we applied corruption to the first $50\%$ of images in each image group while keeping the remaining $50\%$ of images clean.
Baseline methods. We selected five comparison baseline methods that have demonstrated excellent performance in co-saliency object detection, including GCAGC [47], GICD [49], GCoNet [9], DCFM [40] and CADC [48].
[…]
5.2 Comparison on Clean Datasets
We compare our ConceptCoSOD method with baselines on Cosal2015, CoSOD3k, and CoCA datasets. As shown in Table 1, the columns represent the baseline models and our method. The top two values for each metric are highlighted, with red indicating rank-1 and yellow indicating rank-2. For the Cosal2015 dataset, our method achieves five top-2 rankings. On the CoSOD3k dataset, our method secures six top-2 rankings. For the CoCA dataset, our method attains rank-1 across all six metrics. Overall, our method demonstrates the best performance across all three datasets.
@@ 5.2 Comparison on Clean Datasets
GCAGC 0.8029 0.7162 0.0686 0.8139 0.8652 0.8137
GICD 0.8382 0.7484 0.0744 0.8261 0.8774 0.8314
GCoNet 0.8272 0.7472 0.0708 0.8327 0.8808 0.8324
DCFM 0.8486 0.7384 0.0678 0.8446 0.8883 0.8291
CADC 0.8645 0.7723 0.0562 0.8452 0.8943 0.8467
ConceptCoSOD 0.8883 0.7701 0.0534 0.8412 0.9041 0.8467
@@
GCAGC 0.7026 0.6260 0.0797 0.7362 0.8180 0.7674
GICD 0.7361 0.6564 0.0883 0.7461 0.8326 0.7779
GCoNet 0.7322 0.6584 0.0773 0.7569 0.8464 0.7847
DCFM 0.7677 0.6706 0.0712 0.7853 0.8645 0.7949
CADC 0.7457 0.6737 0.0816 0.7559 0.8414 0.7889
ConceptCoSOD 0.8000 0.6981 0.0718 0.7837 0.8664 0.8051
@@
GCAGC 0.3915 0.3766 0.0978 0.4724 0.6878 0.6360
GICD 0.3683 0.4031 0.1455 0.4757 0.6778 0.6311
GCoNet 0.3984 0.4206 0.1147 0.5089 0.7306 0.6517
DCFM 0.4826 0.4697 0.0901 0.5697 0.7646 0.6914
CADC 0.4239 0.4406 0.1275 0.5034 0.6970 0.6569
ConceptCoSOD 0.5652 0.5290 0.0886 0.6091 0.7698 0.7201
5.3 Comparison on Corrupted Datasets
The robustness of our co-salient object detection method is essential, as real-world images often include various corruptions. To assess robustness, we applied common corruptions and an adversarial attack, as presented in Table 2. Specifically, we evaluated a weather-related corruption (frost), common distortions (motion blur, defocus blur, Gaussian noise) [16], and an adversarial attack (Jadena) [14]. The top two scores for each metric are highlighted, with red indicating the highest and yellow indicating the second highest. Results show that our method achieves top performance across all investigated corruption scenarios, including frost, motion blur, defocus blur, Gaussian noise, and adversarial attacks. Our method demonstrates superior robustness across a variety of corruptions.
5.4 Visualization
As shown in Fig. 6, we demonstrate the visualization result comparison between our method and baselines under corruptions. There are three cases, each containing five types of corruption. In most instances, only our ConceptCoSOD method effectively segments the target object, while other methods largely fail to identify it. This outcome demonstrates the superior robustness of our method against corruption compared to others.
Table 2: Co-saliency detection performance on corrupted datasets.
Frost
[…]
GCAGC 0.6714 0.6114 0.0984 0.7233 0.7938 0.7489
GICD 0.7315 0.6589 0.1008 0.7581 0.8256 0.7745
GCoNet 0.6823 0.6196 0.1052 0.7380 0.8049 0.7539
DCFM 0.6848 0.6093 0.1037 0.7462 0.8056 0.7482
CADC 0.7429 0.6683 0.0841 0.7698 0.8275 0.7830
ConceptCoSOD 0.8198 0.7130 0.0809 0.7929 0.8567 0.8055
Motion Blur
[…]
GCAGC 0.6853 0.6113 0.1063 0.7160 0.7935 0.7409
GICD 0.7052 0.6332 0.1088 0.7290 0.8117 0.7567
GCoNet 0.6203 0.5678 0.1105 0.6935 0.7683 0.7247
DCFM 0.6655 0.5948 0.0997 0.7287 0.7969 0.7404
CADC 0.7260 0.6504 0.0907 0.7431 0.8212 0.7686
ConceptCoSOD 0.8178 0.6943 0.0824 0.7718 0.8519 0.7947
Defocus Blur
[…]
GCAGC 0.7057 0.6269 0.1090 0.7317 0.8039 0.7483
GICD 0.7126 0.6310 0.1059 0.7317 0.8090 0.7571
GCoNet 0.5781 0.5320 0.1180 0.6578 0.7320 0.7038
DCFM 0.6684 0.5958 0.0995 0.7352 0.7968 0.7422
CADC 0.6928 0.6162 0.0940 0.7121 0.7913 0.7490
ConceptCoSOD 0.7786 0.6689 0.0888 0.7461 0.8313 0.7761
Jadena
[…]
GCAGC 0.5980 0.5501 0.1256 0.6502 0.7336 0.7025
GICD 0.6794 0.6162 0.1110 0.7111 0.7905 0.7481
GCoNet 0.5627 0.5257 0.1178 0.6447 0.7218 0.7014
DCFM 0.5995 0.5404 0.1097 0.6678 0.7412 0.7112
CADC 0.7176 0.6417 0.0866 0.7380 0.8027 0.7688
ConceptCoSOD 0.8049 0.6996 0.0883 0.7801 0.8465 0.7950
Gaussian Noise
[…]
GCAGC 0.5320 0.5156 0.2074 0.6090 0.6666 0.6383
GICD 0.6853 0.6288 0.1224 0.7200 0.7995 0.7470
GCoNet 0.6054 0.5571 0.1183 0.6887 0.7663 0.7155
DCFM 0.6312 0.5659 0.1112 0.7044 0.7760 0.7217
CADC 0.7141 0.6457 0.0949 0.7461 0.8149 0.7652
ConceptCoSOD 0.7821 0.6857 0.0912 0.7666 0.8340 0.7873
5.5 Ablation Study
Ratio of timestep resampling. Given the validation in Sec. 4.2 that performing denser sampling on the timesteps within the middle intervals is necessary, we further investigate the impact of different sampling ratios on the experimental results. Specifically, we examine the ratios of 1.25, 1.5, 1.75, and 2.0, and compare them with the base version (i.e., performing uniform sampling in all timesteps). As shown in Table 3, all the validated ratio settings outperform the base version in performance as a whole. Among them, the 1.5 ratio yields the best results in the all six metrics. Therefore, we select 1.5 as the default resampling ratio for our experiments.
Table 3: Results of our ConceptCoSOD method under different ratios of timestep resampling.
Cosal2015
[…]
base 0.8605 0.7507 0.0576 0.8236 0.8870 0.8358
ratio 1.25 0.8694 0.7521 0.0593 0.8263 0.8897 0.8353
ratio 1.5 0.8883 0.7701 0.0534 0.8412 0.9041 0.8467
ratio 1.75 0.8689 0.7518 0.0609 0.8235 0.8887 0.8337
ratio 2.0 0.8724 0.7585 0.0563 0.8317 0.8927 0.8406
@@ 5.5 Ablation Study
$E^{max}_{\xi}$$\uparrow$ 0.7331 0.7677 0.7698 0.7536 0.7234 0.6583
$S_{m}$$\uparrow$ 0.6753 0.7053 0.7201 0.7178 0.7018 0.6616
@@ 5.6 Limitation
Our method performs well in general but remains sensitive to challenging cases. We plan to address this by incorporating advanced data augmentation strategies.
5.7 Border Impact
Our method enhances co-salient object detection by integrating semantic guidance, improving robustness in tasks like autonomous driving, medical imaging, and video analysis. This promotes more consistent and interpretable visual recognition in complex scenes. However, such improvements could be misused for surveillance or amplify biases from pre-trained models. We encourage ethical deployment, including fairness evaluation and human oversight, to ensure responsible use.
@@ 6 Conclusion
This paper presents a novel approach, ConceptCoSOD, which leverages text-based semantic guidance for co-salient object detection (Co-SOD). Unlike traditional Co-SOD methods that rely solely on image-based features, ConceptCoSOD introduces a robust (image-text)-to-image framework, enhancing object segmentation accuracy by incorporating semantic information from text. This approach showcases the potential of integrating text-driven semantics into Co-SOD tasks, providing a new direction for future research in vision-language models for complex segmentation tasks. In future work, we aim to enhance the robustness of concept learning to improve the method’s practicality.